OpenStudy (anonymous):

FREE MEDAL>>>While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

HanAkoSolo (jamierox4ev3r):

no the answer would not be the same b/c he changed the opperation from multiplication to addition. If the same number is being multiplied more than once, than you would use an exponent x3^3

OpenStudy (austinl):

\(x^3 \times x^3 \times x^3 \times x^3=x^{12}\)